OpenStudy (anonymous):

If the principal is $200 and the interest rate is 4 percent, what is the simple interest earned in one year?

OpenStudy (error1603):

umm

OpenStudy (anonymous):

What 200$ is it for YOU>:?

OpenStudy (error1603):

If the principal is $200 and the interest rate is 4 percent, what is the simple interest earned in one year? I=Prt=200*.04*1=8. The simple interest earned in one year is?

OpenStudy (anonymous):

you said it was 200$:( you was saying a math

OpenStudy (error1603):

Hello @Rose_Marie2016

OpenStudy (error1603):

$8

OpenStudy (anonymous):

that's hair stuff you name you name hair stuff from rose marie

OpenStudy (error1603):

AFK?

OpenStudy (anonymous):

??????????

OpenStudy (anonymous):

What is AFK

OpenStudy (anonymous):

please do not spam a question with banter

OpenStudy (anonymous):

away from keyboard = AFK

OpenStudy (anonymous):

someone help me im lost

OpenStudy (anonymous):

oK I'M NOT:)

OpenStudy (error1603):

What? Who's spamming?

OpenStudy (anonymous):

Going to slam

OpenStudy (anonymous):

i'm not

OpenStudy (anonymous):

is someone going to help me?

OpenStudy (error1603):

It's $8

OpenStudy (anonymous):

thank you @Error1603

OpenStudy (error1603):

NP

OpenStudy (anonymous):

will you help with my next one?

OpenStudy (anonymous):

@Error1603

OpenStudy (error1603):

Yes

OpenStudy (anonymous):

Sara borrowed $500 for four years at 3 percent interest, compounded annually. Use the formula to calculate the total amount she will have paid back.

OpenStudy (error1603):

The amount she will have to pay back is $562.75. M=P(1+i)^n M=500(1.03)^4 M=500(1.12550881) M=562.75

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

Ryan has an eight-year loan for $6,000. He is being charged an interest rate of 5 percent, compounded annually. Calculate the total amount that he will pay........

OpenStudy (error1603):

He'll pay back is $8,864.73

OpenStudy (anonymous):

What is the total amount that First Consumer Bank will receive after lending Jane $7,000 for three years at an interest rate of 5 percent, compounded annually? .......

OpenStudy (error1603):

Compounded annually will be $1103.38 in interest, for a total of $8,103,38

OpenStudy (anonymous):

thanks for all the help! have a great day

OpenStudy (error1603):

No problem and you too